OpenStudy (priyar):

n identical waves each of intensity Io interfere with each other.The ratio of maximum intensities if the interference is i) coherent ii) incoherent is:

OpenStudy (priyar):

what does coherent interference mean?

OpenStudy (samigupta8):

Generation from a same source

OpenStudy (priyar):

yeah but how will that affect the interference?

OpenStudy (samigupta8):

There we can get constant phase difference all the time... In case of coherent sources

OpenStudy (priyar):

oh.. can you explain how to solve this Q.. coz i still didn't get a clear picture..

OpenStudy (priyar):

@agent0smith

OpenStudy (priyar):

@samigupta8 @ganeshie8

ganeshie8 (ganeshie8):

Is the answer \(n^2\) for coherent and \(n\) for incoherent ?

OpenStudy (priyar):

just the ratio is given which is correct... so your answer is correct1 now, pls tell me hoe to solve..

OpenStudy (priyar):

the ratio is n

OpenStudy (priyar):

@ganeshie8 ?

ganeshie8 (ganeshie8):

Consider the electric field of one light wave at a specific point, say, \(x = 0\) : \[E_1 = E_0\sin(\omega t)\tag{1}\]

ganeshie8 (ganeshie8):

If the light waves are coherent (same frequency, phase), then the electric fields of remaining waves also follow the same equation as above.

ganeshie8 (ganeshie8):

Add up the electric fields and get \[E_{net} = nE_0 \sin(\omega t)\]

ganeshie8 (ganeshie8):

Next, use the fact that \(I \propto E^2 \)

OpenStudy (priyar):

ok.. so max. intensity will be n^2 .. but for the second case we can't add them as they won't have const. phase diff right?

OpenStudy (priyar):

@ganeshie8

OpenStudy (priyar):

@mayankdevnani

OpenStudy (priyar):

@ganeshie8 how did you get n as max. intensity for incoherent?

OpenStudy (mayankdevnani):

i got `n` as answer ?

OpenStudy (mayankdevnani):

hope this is correct !

OpenStudy (priyar):

thats right! Now please explain

OpenStudy (mayankdevnani):

yipee !

OpenStudy (mayankdevnani):

well @ganeshie8 is correct ! but i have a different reasoning

OpenStudy (priyar):

oh..what is it?

OpenStudy (mayankdevnani):

We know that, Interference of two wave can be written as :- \[\large \bf I=I_1+I_2+2\sqrt{I_1I_2}\cos \phi \]

OpenStudy (priyar):

yes

OpenStudy (mayankdevnani):

Case 1 :- When coherent since for coherent source,`phi` doesn't matter at all. For max intensity, cos phi=1 So, \[\large \bf I_{\max}=I_1+I_2+2\sqrt{I_1I_2}=(\sqrt I_1 +\sqrt I_2)^2\] For `n waves`, \[\large \bf I_{\max}=(\sqrt I_1+\sqrt I_2 + .....)^2=(n \sqrt I_0)^2=n^2 I_0\]

OpenStudy (mayankdevnani):

similarly apply for incoherent source

OpenStudy (priyar):

the second case is what i am doubtful in..

OpenStudy (mayankdevnani):

okay

OpenStudy (mayankdevnani):

Case 2 :- Incoherent Source since `phi` depends. So for `n waves` we take average , \[\large \bf <\cos \phi>=0\] \[\large \bf I=I_1+I_2\] for `n waves`, \[\large \bf I=I_1+I_2+.........=n I_0\]

OpenStudy (mayankdevnani):

now take ratio,you'll get `n`

OpenStudy (priyar):

oh nice! Thanks!

OpenStudy (mayankdevnani):

no problem :)

OpenStudy (mayankdevnani):

THUMBS UP :D @ganeshie8

ganeshie8 (ganeshie8):

Nice!

OpenStudy (mayankdevnani):

<3

OpenStudy (priyar):

Thanks all!

OpenStudy (mayankdevnani):

yw :)